OpenStudy (anonymous):

Solve the given linear system -2x+3y=8 3x+5y=-12

OpenStudy (anonymous):

With a graphing calculator, substitution, elimination, or matrix. Graphing calculator is always the easiest.

OpenStudy (anonymous):

Put the equations in slope-intercept form and graph it on your calculator. Then find the intercept, you hit the second key and then trace, then hit 5. Click enter 3 times since your lines are linear, you do not need to worry about 2 intersections. Then your x and y value should show. Good luck my friend.

OpenStudy (anonymous):

but we dont have graphing calculator

OpenStudy (anonymous):

Here's an online graphing calculator, https://sites.google.com/site/northwesternhighschoolptsa/free-online-graphing-calculator

OpenStudy (anonymous):

It's free, and if your teacher doesn't let you use that then use the elimination method.

OpenStudy (anonymous):

thanks.. does it provide solutions?

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

Solve for y in your equations and then plug it in to the calculator.

OpenStudy (anonymous):

I could get you started on the elimination method.

OpenStudy (anonymous):

-2x+3y=8 3x+5y=-12 You want to cancel out one of the variables, -2x and 3x looks like the easiest to me. Make sure one of them is negative and one of them is positive so you can cancel out both of them. 3{-2x+3y=8} 2{3x+5y=-12} (Distribute 3 and 2 to everything, including that 8 and -12) -6x-9y=-24 6x+10y=-24 (x cancels out, 6-6=0) y=-48 Looks like you got an easy one. Plug in -48 into any of that equations for y. (I'm just going to use the first one.) -2x+3(-48)=8 -2x-144=8 +144 -2x=152 x=-76 Answer: x=-76 y=-48 Or, (-76,-48)

OpenStudy (anonymous):

where did you get 3&2?

OpenStudy (anonymous):

and how did you get the 48? im really sorry if i dont get it much

OpenStudy (anonymous):

i think i get it..

OpenStudy (anonymous):

now the second one is x=-84

OpenStudy (anonymous):

The 3 and 2 was something I had to come up on my own. You can multiply 3 and 2 to get a 6. Only one has to be a negative so when you add the equations together, a positive plus a negative will cancel out because they are the same number. I got a -48 by adding -24 and -24